OpenStudy (anonymous):

Landing with a speed of 81.9 m/s, and traveling due south, a jet comes to rest in 949m. Assuming the jet slows with constant acceleration, find the magnitude and direction of its acceleration.

OpenStudy (anonymous):

Which formula should I use?

OpenStudy (goformit100):

use Vector Formula.

OpenStudy (anonymous):

that only tells me A + B -> AB (Resultant)..

OpenStudy (anonymous):

I know it would have to be accelerating "due north" at a certain (m/s2).

OpenStudy (anonymous):

How do I solve for acceleration please?

OpenStudy (anonymous):

\[v=v_0-at\] when it comes to rest:\[v=0 \rightarrow t=\frac{ v_0 }{ a }\]distance is calculated like:\[S=v_0t-\frac{ 1 }{ 2}at^2=v_0\left( \frac{ v_0 }{ a} \right)-\frac{ 1 }{ 2 }a\left( \frac{ v_0 }{ a} \right)^2=\frac{ v_0^2 }{ 2a }\rightarrow a=\frac{ v_0^2 }{ 2S }\] You know Vo=81.9 m/sec and S=949

OpenStudy (anonymous):

Thank you, I used \[\frac{ v _{0}^{2} }{ 2S }\] v initial = 81.9 m/sec and S=949m to get 3.53 m/sec^2 Can you please go over the derivation again to eliminate t? I do not fully understand the last step.

OpenStudy (anonymous):

Just replace the expresion of time to come to rest t=Vo/a in the expression of distance run in this time: S=Vo·t-1/2a·t^2